# Mobile-Bench-v2: A More Realistic and Comprehensive Benchmark for VLM-based Mobile Agents

### v2

@@ 1 Introduction
LLM-based mobile agents (wang2023enabling, ; ding2024mobileagent, ) are increasingly popular due to their capability to interact directly with mobile Graphical User Interfaces (GUIs) and their potential to manage daily tasks autonomously. Unfortunately, LLM-based agents cannot well comprehend the mobile GUI structure and widget functionality, base on text such as Visual-Hierarchical, XML, HTML, or Accessible-Visited Trees. Recent studies (ma2024coco, ; zhang2024mobileexperts, ) indicate VLMs can provide a more comprehensive understanding of GUIs. This has led to mobile benchmark replacing foundational models with VLMs, resulting in benchmarks for end-to-end mobile tasks based on GUI pages (wang2024mobileagentbench, ; xu2024androidlab, ; rawles2024androidworld, ).
@@
PixelHelp (li2020mapping, ) 187 4.2 $\sim$800 Single
MoTIF (burns2021mobile, ) 480 4.5 $\sim$21K Single
AitW (rawles2024androidinthewild, ) 1,539 6.5 $\sim$510K Single
AitZ (zhang2024android, ) 506 7.5 $\sim$18K Single
AMEX (chai2024amex, ) 341 12.8 $\sim$104K Single
ScreenSpot cheng2024seeclick $\sim$1,200 1 $\sim$600 Dot
MobileAIBench (murthy2024mobileaibench, ) * * * Dot *
AgentBench (wang2024mobileagentbench, ) 100 20 $\sim$2k Multiple
GUI Odeyssey (lu2024gui, ) 7,735 15.4 * Single
Mobile-Bench (deng2024mobile, ) 832 * 14,144 Multiple
Mobile-Env (zhang2023mobile, ) 150 * * Multiple
AndroidLab (xu2024androidlab, ) 10.5k 8.98 94.3k Multiple
AndroidWorld (rawles2024androidworld, ) 116 * * Multiple
@@
Existing VLM-based Agent benchmarks can be divided into two categories: (1) Online evaluation involves the agent executing operations on a real device based on the user’s high-level instructions. These benchmarks directly determine the success rate by checking the widget values in the final GUI and allow agents to complete tasks through multi-path. However, due to the instability of the device environment, such as OS updates, APP updates, and user preference records, online benchmarks (murthy2024mobileaibench, ; deng2024mobile, ; wang2024mobileagentbench, ) step rewards are fluctuating and unstable. Meanwhile, a considerable portion of task evaluation relies on the process rather than solely on the outcome. (2) Offline evaluation uses static datasets where the golden path is pre-executed on the device, with actions and screenshots saved offline. The agent generates the current action based on each step’s GUI, instructions and action history, which action trajectory is formulated as single-path. Although offline benchmarks (chai2024amex, ; cheng2024seeclick, ; rawles2024androidinthewild, ) are more convenient, considering the diversity of agent task solutions, the agent’s good performance may only represent a good fit to the preferences encoded in the current benchmark annotations, but does not necessarily indicate robustness or the ability to handle multi-path solutions. More critically, benchmarks such as MobileAgentBench (wang2024mobileagentbench, ) and AutoDroid (wen2024autodroid, ) are constructed on real devices and evaluated within Google apps using the Android Accessibility Service; these apps feature overly clean pages without task-irrelevant ads, buttons, and pop-ups. In real-world scenarios, users may not be able to provide such precise and full instructions all at once (wang2024learning, ). Overall, existing benchmarks have following limitations, including a lack of multi-path evaluation, overly clean testing environments, and overly explicit instructions.
@@
To address the above problems, we form a new benchmark named Mobile-Bench-v2 based on Mobile-Bench (deng2024mobile, ). Mobile-Bench-v2 has the following improvements: (1) Instruction Generation Method: Based on the random walk graph-structured corpus of Mobile3M (wu2024mobilevlm, ), we use the GIAS (Generating Instructions From GUI Action Sequences) to construct 12k instructions. Each instruction is generated via task templates and slot information based on the corresponding trajectories. Slot information is derived from keynode actions in the execution trajectory, which can be used as stable step reward signals during the evaluation stage. (2) Offline Multi-path Evaluation: Combining the advantages of both online and offline environments, we propose a multi-path evaluation approach. We allow the agent to execute in a single-path manner and compare the result with the golden path. Alternatively, the agent is also allowed to perform action search within the graph corpus and accumulate step rewards, as it does during online evaluation. Both approaches are jointly adopted to comprehensively evaluate the agent’s performance. (3) Realistic Noisy Environment: We collect an additional sub-dataset named Mobile-Bench-Noisy to build realistic noisy evaluation. Several apps with substantial ads and pop-ups are specifically selected to collect this type of trajectory. Additionally, we contaminate AITZ by inserting ads into original normal trajectories to build AITZ-Noise. (4) Active Interactive Evaluation: We construct a sub-dataset named Mobile-Bench-Ambiguous, which allows agents to ask when necessary during task execution. Full instructions are pre-constructed and then progressively simplified into ambiguous instructions through slot-based extraction. Questions and answers are built based on slot information and then assigned to the corresponding GUI. Since none of the existing agent frameworks incorporates a dedicated questioning stage, we allow the agent to decide whether to raise a question before making a decision. Questions related to action decisions or UI element functionalities are rejected to ensure the evaluation’s fairness.
@@ 2.1 Mobile Agents
Large language models (achiam2023gpt, ) emerge as autonomous agents (lieffects, ; wen2023empowering, ) in the mobile domain and garner considerable attention. With the rapid development of vision-language models (VLMs), multimodal researchers build mobile GUI agents (yang2023appagent, ; zheng2024gpt, ) and multi-agent frameworks (ding2024mobileagent, ; li2024appagent, ; wang2024mobile, ) based on closed-source VLMs. Meanwhile, some researchers focus on training agents with stronger element grounding (cheng2024seeclick, ; hong2024cogagent, ; wu2024atlas, ), page navigation (niu2024screenagent, ; lu2024gui, ; gou2024navigating, ), GUI understanding (chai2024amex, ; you2024ferret, ; baechler2024screenai, ) and task planning capabilities (zhang2024llamatouch, ; nong2024mobileflow, ; xu2024aguvis, ) based on open-source VLMs. In addition, Digirl (bai2024digirl, ) and Distrl (wang2024distrl, ) uses joint online and offline reinforcement learning to enhance the generalization of mobile agents and mitigate performance degradation when facing APP updates and unseen APPs. Some researchers (qinghong2024showui, ) explore optimizing VLM structures. For example, Dorka (dorka2024training, ) optimizes the encoder by incorporating historical images and actions as input.
@@ 2.2 Mobile Agent Benchmarks
As shown in Table 1, AndroidEnv (toyama2021androidenv, ) and MobilEnv (zhang2023mobile, ) are the first to create LLM agent evaluation environments based on reinforcement learning. Mobile-Bench (deng2024mobile, ) and AppBench (wang2024appbench, ) introduce online benchmarks combining API and GUI, while MobileAgentBench (wang2024mobileagentbench, ) establishes the first fully automated multimodal benchmark for VLM-based GUI agents. More offline benchmarks (li2020mapping, ; burns2021mobile, ; murthy2024mobileaibench, ) are released, which are primarily categorized into GUI understanding and task-oriented. (1) For task-oriented benchmarks, AITW (rawles2024androidinthewild, ) and AITZ (zhang2024android, ) create large-scale benchmarks based on Google apps, while AMEX (chai2024amex, ) supplements these benchmarks by adding data for GUI understanding with similar app types. ScreenSpot (cheng2024seeclick, ), Mobile3M (wu2024mobilevlm, ), and GUIOdyssey (lu2024gui, ) focus on more granular element grounding and task planning. (2) Rico (deka2017rico, ) is the first non-annotated GUI corpus, followed by ScreenQA (hsiao2022screenqa, ), Widget Caption (li2020widget, ), and Screen2words (wang2021screen2words, ), which is for Q&A, widget understanding, and page summarization. Subsequently, Mind2web (deng2024mind2web, ) incorporates additional GUI data of varying sizes, and Meta-GUI (sun2022meta, ) provides tasks for multi-round dialogues.
@@ 3.1 Mobile Task Formulation
For mobile agents, there are four essential capabilities: (i) Overall Planning to determine the action step sequences. (ii) Action Thought to produce an action description at each step (e.g., “open the flight detail page”), (iii) Element Grounding to identify a widget (e.g., “[Click]($x_{1}$,$y_{1}$)”) on the GUI, (iiii) Action Reflection to determine whether the next GUI matches the expectation. Given a mobile screenshot $\mathcal{S}$ (e.g., a Ctrip screenshot on Android) and a task $\mathcal{T}$ (e.g., “Book a flight ticket from Chengdu to Beijing Sep.15 for me.”), a GUI agent should generate a sequence of executable actions. Specifically, at time step $t$, the agent should select an action $\bm{a}_{t}$ from the action space $\mathcal{A}$, which includes three types of actions: (1) Click. (2) Scroll. (3) Input.
@@
Based on the current environment observation $\mathcal{S}_{t}$, the action history $\mathcal{H}_{1:t-1}$=$\{\hat{a}_{1},\hat{a}_{2},...,\hat{a}_{t-1}\}$, and the last step refection $\bm{f}_{t-1}$, the GUI agent will generate plan $\mathcal{P}_{t}$:
@@ 3.2 Data Construction
Generating common instructions from action trajectories. When constructing the Mobile3M graph corpus, the key challenge is annotating instructions for each trajectory that closely align with the intended actions. Building on Netscape (murty2024nnetscape, )’s fine-tuning of web agents to eliminate redundant actions from action sequences, two key points for pairing trajectories and instructions are the Intent Understanding and the Slot Matching between different GUIs: (i) Using intents behind actions instead of themselves can reduce ambiguity because coordinate-based actions without GUI pages cannot accurately reproduce the action scenes. On the one hand, buttons with the same textview may have entirely different meanings on the same GUI. In Figure 7, the same ‘plus’ button represents adding ‘Hazelnut Latte’ and ‘Cookie Mocha’ respectively. On the other hand, UI elements without textual descriptions fail to reflect their underlying functionality. For example, the sorting button in Ctrip is a triangle-shaped graphical button, recording the shape of this button is much less clear than capturing its sorting functionality. (ii) Filling predefined instruction templates with slot information can serve as a reward signal at key nodes. GUI agent tasks are inherently multi-path in nature. Single-path annotations with unstable preference may lead to a sharp performance drop on unseen tasks and further compromise the accuracy of performance evaluation. Slot filling allows a single instruction template to match multiple distinct trajectories, as long as they share the same key nodes. This also serves as the foundation for multi-path evaluation. The details of Mobile3m are described in Appendix A, and data distribution is shown in Figure 5.
Building on the above findings, we propose a slot-based instruction annotation method named GIAS (Generating Instructions From Mobile UI Action Sequences), which is shown in Figure 1. The whole process is as follows: (1) multi-path sampling based on fixed start and end GUIs; (2) GUI content annotation; (3) action intent inference; (4) extracting slot information from GUI changes; (5) filling instruction templates with the slot; (6) deduplication and simplification. The entire process is explained in detail in Algorithm 1. Specifically, we choose paths that start from nodes with the same name in Mobile3M and end at homogeneous nodes with different names (Homogeneous nodes refer to pages whose similarity or the number of identical UI elements exceeds the threshold (lu2006field, )). Considering its diversity, each trajectory includes at least two different types of actions and minimises the proportion of intermediate homogeneous pages. Throughout the entire annotation process, only the verification step involves a closed-source model, while all other steps are performed by open-source models without any human intervention. More details on GIAS are in Appendix B.
@@
1:Start Page, $P_{0}$; End Page, $P_{t}$; Trajectory $\sigma$; Page Description $\mathcal{D}$; GUI Pages, $s$; Action, $a$; Action Intent, $T$; Page Slot, $C$; Instruction, $I$; Task, $\mathcal{T}$;
@@
6:         $\mathcal{D}_{s_{ij}}\leftarrow\mathrm{VLM}(s_{ij},P_{t})$triangleright Get text descriptions from GUIs
7:         $T_{i_{j:j+1}}\leftarrow\mathcal{D}_{s_{ij}},\mathcal{D}_{s_{i,j+1}},a(s_{ij}\to s_{i,j+1})$triangleright Get indent between two actions
8:         $C_{i_{j:j+1}}\leftarrow\mathcal{D}_{s_{ij}},\mathcal{D}_{s_{i,j+1}}$triangleright Get Page Slot based on pre-settings
9:         $I_{ij}\leftarrow(C_{i:t},T_{i:t})\sim\text{Uniform}(\mathcal{\gamma})$triangleright Fill Templates with the slot
@@
14:         Discard $I_{j}$ from $\mathcal{I}$triangleright Make sure there are no highly similar instructions
@@
17:$\mathcal{T},Q\leftarrow\mathrm{Veri}\ \{\mathcal{I},P,F_{S}\}$triangleright for all Trajectories $\sigma_{i}$, ensuring no redundant steps
@@
Noisy app and ambiguous instruction data. Mobile-Bench-Noisy is primarily derived from manual annotation and contamination in existing data: (1) For manual annotation, we select apps from third-party markets; these apps contain unavoidable ads and pop-ups. When performing actions on these apps, we do not handle the following scenarios in advance: login, update, permission settings, ad pop-ups, and VIP subscriptions. In some instructions, to test the agent’s response in unexpected situations, we deliberately click on ad pages incorrectly to see if it can recover from divergent paths. All noisy steps and their redirect pages are additionally marked, making this subset adaptable for agents with a rollback mechanism as well. (2) For data contamination, we randomly insert at least one ad (randomly collected from the Google Store app) into AITZ (zhang2024android, ) trajectories, which is a high-quality subset of AITW (rawles2024androidinthewild, ). For Mobile-Bench-Ambiguous, we first construct the full instructions, annotate action trajectories, and remove slot information to build ambiguous instructions. Multiple sets of interactive Q&A are assigned to corresponding GUIs. For example, as shown in figure 2, the full instruction is: ‘I want a 16GB + 512GB MacBook Pro M4 in the Midnight version.’, while the ambiguous instruction is: ‘I want to buy a MacBook.’. Information such as “16GB + 512GB”, “Pro M4”, and “Midnight version” are treated as three slots, each assigned to the corresponding GUI for step reward. More details can be seen in the appendix C.3.

### v3

@@ 1 Introduction
LLM-based mobile agents (Wang et al., 2023; Ding, 2024) are increasingly popular due to their capability to interact directly with mobile Graphical User Interfaces (GUIs) and their potential to manage daily tasks autonomously. Unfortunately, LLM-based agents cannot well comprehend the mobile GUI structure and widget functionality, base on text such as Visual-Hierarchical, XML, HTML, or Accessible-Visited Trees. Recent studies (Ma et al., 2024; Zhang et al., 2024a) indicate VLMs can provide a more comprehensive understanding of GUIs. This has led to mobile benchmark replacing foundational models with VLMs, resulting in benchmarks for end-to-end mobile tasks based on GUI pages (Wang et al., 2024c; Xu et al., 2024a; Rawles et al., 2024).
@@
PixelHelp (Li et al., 2020a) 187 4.2 $\sim$800 Single
MoTIF (Burns et al., 2021) 480 4.5 $\sim$21K Single
AitW (Rawles et al., 2023) 1,539 6.5 $\sim$510K Single
AitZ (Zhang et al., 2024b) 506 7.5 $\sim$18K Single
AMEX (Chai et al., 2024) 341 12.8 $\sim$104K Single
ScreenSpot Cheng et al. (2024) $\sim$1,200 1 $\sim$600 Dot
MobileAIBench (Murthy et al., 2024) * * * Dot *
AgentBench (Wang et al., 2024c) 100 20 $\sim$2k Multiple
GUI Odeyssey (Lu et al., 2024) 7,735 15.4 * Single
Mobile-Bench (Deng et al., 2024) 832 * 14,144 Multiple
Mobile-Env (Zhang et al., 2023) 150 * * Multiple
AndroidLab (Xu et al., 2024a) 10.5k 8.98 94.3k Multiple
AndroidWorld (Rawles et al., 2024) 116 * * Multiple
@@
Existing VLM-based Agent benchmarks can be divided into two categories: (1) Online evaluation involves the agent executing operations on a real device based on the user’s high-level instructions. These benchmarks directly determine the success rate by checking the widget values in the final GUI and allow agents to complete tasks through multi-path. However, due to the instability of the device environment, such as OS updates, APP updates, and user preference records, online benchmarks (Murthy et al., 2024; Deng et al., 2024; Wang et al., 2024c) step rewards are fluctuating and unstable. Meanwhile, a considerable portion of task evaluation relies on the process rather than solely on the outcome. (2) Offline evaluation uses static datasets where the golden path is pre-executed on the device, with actions and screenshots saved offline. The agent generates the current action based on each step’s GUI, instructions and action history, which action trajectory is formulated as single-path. Although offline benchmarks (Chai et al., 2024; Cheng et al., 2024; Rawles et al., 2023) are more convenient, considering the diversity of agent task solutions, the agent’s good performance may only represent a good fit to the preferences encoded in the current benchmark annotations, but does not necessarily indicate robustness or the ability to handle multi-path solutions. More critically, benchmarks such as MobileAgentBench (Wang et al., 2024c) and AutoDroid (Wen et al., 2024) are constructed on real devices and evaluated within Google apps using the Android Accessibility Service; these apps feature overly clean pages without task-irrelevant ads, buttons, and pop-ups. In real-world scenarios, users may not be able to provide such precise and full instructions all at once (Wang et al., 2024f). Overall, existing benchmarks have following limitations, including a lack of multi-path evaluation, overly clean testing environments, and overly explicit instructions.
@@
To address the above problems, we form a new benchmark named Mobile-Bench-v2 based on Mobile-Bench (Deng et al., 2024). Mobile-Bench-v2 has the following improvements: (1) Instruction Generation Method: Based on the random walk graph-structured corpus of Mobile3M (Wu et al., 2024a), we use the GIAS (Generating Instructions From GUI Action Sequences) to construct 12k instructions. Each instruction is generated via task templates and slot information based on the corresponding trajectories. Slot information is derived from keynode actions in the execution trajectory, which can be used as stable step reward signals during the evaluation stage. (2) Offline Multi-path Evaluation: Combining the advantages of both online and offline environments, we propose a multi-path evaluation approach. We allow the agent to execute in a single-path manner and compare the result with the golden path. Alternatively, the agent is also allowed to perform action search within the graph corpus and accumulate step rewards, as it does during online evaluation. Both approaches are jointly adopted to comprehensively evaluate the agent’s performance. (3) Realistic Noisy Environment: We collect an additional sub-dataset named Mobile-Bench-Noisy to build realistic noisy evaluation. Several apps with substantial ads and pop-ups are specifically selected to collect this type of trajectory. Additionally, we contaminate AITZ by inserting ads into original normal trajectories to build AITZ-Noise. (4) Active Interactive Evaluation: We construct a sub-dataset named Mobile-Bench-Ambiguous, which allows agents to ask when necessary during task execution. Full instructions are pre-constructed and then progressively simplified into ambiguous instructions through slot-based extraction. Questions and answers are built based on slot information and then assigned to the corresponding GUI. Since none of the existing agent frameworks incorporates a dedicated questioning stage, we allow the agent to decide whether to raise a question before making a decision. Questions related to action decisions or UI element functionalities are rejected to ensure the evaluation’s fairness.
@@ 2.1 Mobile Agents
Large language models (Achiam et al., 2023) emerge as autonomous agents (Li et al., 2024b; Wen et al., 2023) in the mobile domain and garner considerable attention. With the rapid development of vision-language models (VLMs), multimodal researchers build mobile GUI agents (Yang et al., 2023; Zheng et al., 2024) and multi-agent frameworks (Ding, 2024; Li et al., 2024c; Wang et al., 2024b) based on closed-source VLMs. Meanwhile, some researchers focus on training agents with stronger element grounding (Cheng et al., 2024; Hong et al., 2024; Wu et al., 2024b), page navigation (Niu et al., 2024; Lu et al., 2024; Gou et al., 2024), GUI understanding (Chai et al., 2024; You et al., 2024; Baechler et al., 2024) and task planning capabilities (Zhang et al., 2024c; Nong et al., 2024; Xu et al., 2024b) based on open-source VLMs. In addition, Digirl (Bai et al., 2024) and Distrl (Wang et al., 2024e) uses joint online and offline reinforcement learning to enhance the generalization of mobile agents and mitigate performance degradation when facing APP updates and unseen APPs. Some researchers (Qinghong Lin et al., 2024) explore optimizing VLM structures. For example, Dorka (Dorka et al., 2024) optimizes the encoder by incorporating historical images and actions as input.
@@ 2.2 Mobile Agent Benchmarks
As shown in Table 1, AndroidEnv (Toyama et al., 2021) and MobilEnv (Zhang et al., 2023) are the first to create LLM agent evaluation environments based on reinforcement learning. Mobile-Bench (Deng et al., 2024) and AppBench (Wang et al., 2024a) introduce online benchmarks combining API and GUI, while MobileAgentBench (Wang et al., 2024c) establishes the first fully automated multimodal benchmark for VLM-based GUI agents. More offline benchmarks (Li et al., 2020a; Burns et al., 2021; Murthy et al., 2024) are released, which are primarily categorized into GUI understanding and task-oriented. (1) For task-oriented benchmarks, AITW (Rawles et al., 2023) and AITZ (Zhang et al., 2024b) create large-scale benchmarks based on Google apps, while AMEX (Chai et al., 2024) supplements these benchmarks by adding data for GUI understanding with similar app types. ScreenSpot (Cheng et al., 2024), Mobile3M (Wu et al., 2024a), and GUIOdyssey (Lu et al., 2024) focus on more granular element grounding and task planning. (2) Rico (Deka et al., 2017) is the first non-annotated GUI corpus, followed by ScreenQA (Hsiao et al., 2022), Widget Caption (Li et al., 2020b), and Screen2words (Wang et al., 2021), which is for Q&A, widget understanding, and page summarization. Subsequently, Mind2web (Deng et al., 2023) incorporates additional GUI data of varying sizes, and Meta-GUI (Sun et al., 2022) provides tasks for multi-round dialogues.
@@ 3.1 Mobile Task Formulation
For mobile agents, there are four essential capabilities: (i) Overall Planning to determine the action step sequences. (ii) Action Thought to produce an action description at each step (e.g., “open the flight detail page”), (iii) Element Grounding to identify a widget (e.g., “[Click]($x_{1}$$y_{1}$)”) on the GUI, (iiii) Action Reflection to determine whether the next GUI matches the expectation. Given a mobile screenshot $\mathcal{S}$ (e.g., a Ctrip screenshot on Android) and a task $\mathcal{T}$ (e.g., “Book a flight ticket from Chengdu to Beijing Sep.15 for me.”), a GUI agent should generate a sequence of executable actions. Specifically, at time step $t$, the agent should select an action $\bm{a}_{t}$ from the action space $\mathcal{A}$, which includes three types of actions: (1) Click. (2) Scroll. (3) Input.
@@
Based on the current environment observation $\mathcal{S}_{t}$, the action history $\mathcal{H}_{1:t-1}$$\{\hat{a}_{1},\hat{a}_{2},...,\hat{a}_{t-1}\}$, and the last step refection $\bm{f}_{t-1}$, the GUI agent will generate plan $\mathcal{P}_{t}$:
@@ 3.2 Data Construction
Generating common instructions from action trajectories. When constructing the Mobile3M graph corpus, the key challenge is annotating instructions for each trajectory that closely align with the intended actions. Building on Netscape (Murty et al., 2024)’s fine-tuning of web agents to eliminate redundant actions from action sequences, two key points for pairing trajectories and instructions are the Intent Understanding and the Slot Matching between different GUIs: (i) Using intents behind actions instead of themselves can reduce ambiguity because coordinate-based actions without GUI pages cannot accurately reproduce the action scenes. On the one hand, buttons with the same textview may have entirely different meanings on the same GUI. In Figure 7, the same ‘plus’ button represents adding ‘Hazelnut Latte’ and ‘Cookie Mocha’ respectively. On the other hand, UI elements without textual descriptions fail to reflect their underlying functionality. For example, the sorting button in Ctrip is a triangle-shaped graphical button, recording the shape of this button is much less clear than capturing its sorting functionality. (ii) Filling predefined instruction templates with slot information can serve as a reward signal at key nodes. GUI agent tasks are inherently multi-path in nature. Single-path annotations with unstable preference may lead to a sharp performance drop on unseen tasks and further compromise the accuracy of performance evaluation. Slot filling allows a single instruction template to match multiple distinct trajectories, as long as they share the same key nodes. This also serves as the foundation for multi-path evaluation. The details of Mobile3m are described in Appendix A, and data distribution is shown in Figure 5.
Building on the above findings, we propose a slot-based instruction annotation method named GIAS (Generating Instructions From Mobile UI Action Sequences), which is shown in Figure 1. The whole process is as follows: (1) multi-path sampling based on fixed start and end GUIs; (2) GUI content annotation; (3) action intent inference; (4) extracting slot information from GUI changes; (5) filling instruction templates with the slot; (6) deduplication and simplification. The entire process is explained in detail in Algorithm 1. Specifically, we choose paths that start from nodes with the same name in Mobile3M and end at homogeneous nodes with different names (Homogeneous nodes refer to pages whose similarity or the number of identical UI elements exceeds the threshold (Lu et al., 2006)). Considering its diversity, each trajectory includes at least two different types of actions and minimises the proportion of intermediate homogeneous pages. Throughout the entire annotation process, only the verification step involves a closed-source model, while all other steps are performed by open-source models without any human intervention. More details on GIAS are in Appendix B.
@@
1:Start Page, $P_{0}$; End Page, $P_{t}$; Trajectory $\sigma$; Page Description $\mathcal{D}$; GUI Pages, $s$; Action, $a$; Action Intent, $T$; Page Slot, $C$; Instruction, II; Task, $\mathcal{T}$;
@@
6:    $\mathcal{D}_{s_{ij}}\leftarrow\mathrm{VLM}(s_{ij},P_{t})$ $\triangleright$ Get text descriptions from GUIs
7:    $T_{i_{j:j+1}}\leftarrow\mathcal{D}_{s_{ij}},\mathcal{D}_{s_{i,j+1}},a(s_{ij}\to s_{i,j+1})$ $\triangleright$ Get indent between two actions
8:    $C_{i_{j:j+1}}\leftarrow\mathcal{D}_{s_{ij}},\mathcal{D}_{s_{i,j+1}}$ $\triangleright$ Get Page Slot based on pre-settings
9:    $I_{ij}\leftarrow(C_{i:t},T_{i:t})\sim\text{Uniform}(\mathcal{\gamma})$ $\triangleright$ Fill Templates with the slot
@@
14:    Discard $I_{j}$ from $\mathcal{I}$ $\triangleright$ Make sure there are no highly similar instructions
@@
[…]
@@
Noisy app and ambiguous instruction data. Mobile-Bench-Noisy is primarily derived from manual annotation and contamination in existing data: (1) For manual annotation, we select apps from third-party markets; these apps contain unavoidable ads and pop-ups. When performing actions on these apps, we do not handle the following scenarios in advance: login, update, permission settings, ad pop-ups, and VIP subscriptions. In some instructions, to test the agent’s response in unexpected situations, we deliberately click on ad pages incorrectly to see if it can recover from divergent paths. All noisy steps and their redirect pages are additionally marked, making this subset adaptable for agents with a rollback mechanism as well. (2) For data contamination, we randomly insert at least one ad (randomly collected from the Google Store app) into AITZ (Zhang et al., 2024b) trajectories, which is a high-quality subset of AITW (Rawles et al., 2023). For Mobile-Bench-Ambiguous, we first construct the full instructions, annotate action trajectories, and remove slot information to build ambiguous instructions. Multiple sets of interactive Q&A are assigned to corresponding GUIs. For example, as shown in figure 2, the full instruction is: ‘I want a 16GB + 512GB MacBook Pro M4 in the Midnight version.’, while the ambiguous instruction is: ‘I want to buy a MacBook.’. Information such as “16GB + 512GB”, “Pro M4”, and “Midnight version” are treated as three slots, each assigned to the corresponding GUI for step reward. More details can be seen in the appendix C.3.